OpenStudy (anonymous):

If f(x) = 2x + 2, find f(5). Choose one answer. a. 10 b. 12 c. 9 d. 5

OpenStudy (anonymous):

Plug five in for x then solve the equation.

OpenStudy (anonymous):

f(5)=2(5)+2 =10+2 = 12

OpenStudy (the_fizicx99):

Plug in 5 for every x term, f(5) = 2(5) 2 f(5) = 10 + 2

OpenStudy (the_fizicx99):

@larryboxaplenty please stop giving out answer U_U